OpenStudy (anonymous):

x=7-2y y=3x plz help!!!!

OpenStudy (lgbasallote):

since y = 3x you can say that x = 7 - 2(3x) right? (i just substituted 3x into y)

OpenStudy (anonymous):

so it would be x=7-6y?

OpenStudy (lgbasallote):

but it's x already

OpenStudy (lgbasallote):

x = 7 - 6x

OpenStudy (anonymous):

o yea i meant that my bad so wat do i do next?

OpenStudy (lgbasallote):

you solve for x

OpenStudy (lgbasallote):

im out now..hope you figure out the rest

OpenStudy (precal):